math

Write an equation for the function that is described by the given characteristics.
The shape of f(x) = x2, but moved eight units to the left, nine units upward, and then reflected in the x-axis
Answer #1

moving 8 units to the left:

f(x)=(x+8)2

nine units upward:

f(x)=(x+8)2+9

reflected about the x axis:

f(x)(-1)=[(x+8)2+9](-1)

so f(x)(-1)=-9-(x+8)2

answered by: Terence